Debt: 500 4% coupon bonds outstanding, with a par value of $1,000, 10 years to maturity, trading at $950 and making semi-annual payments.

Equity: 20,000 shares outstanding and trading at a price of $40 per share. The current dividend is $1 per share and the expected earnings growth rate is 5%.

A) What is the WACC?  Assume a tax rate of 30%
Answer #1

Debt component:
Par value or face value=$1000
Present value=$950
Coupon rate=4%
As the coupon payment is made semiannually, the semiannual coupon rate=4%/2=0.02
Coupon payment=(Semiannual coupon rate)*(Face value)=0.02*1000=20
Time period=10 years
As the coupon payment is made semiannually, the number of periods=10*2=20
Cost of debt is the yield to maturity that we need to determine.

А 1 Face value 1000 2 Present value -950 3 Coupon payment 20 4 Number of periods 5 Yield to maturity= 2.32% 6 Formula used: R

Semiannual yield to maturity or the cost of debt=2.32%
Annualized yield to maturity=2.32%*2=4.64%

Equity component:
Current dividend=$1
Current price=$40
Growth rate=5%
Now, Current price=(Current dividend)*(1+Growth rate)/(Cost of equity - Growth rate)
40=(1)*(1+5%)/(Cost of equity - 5%)
40=1.05/(Cost of equity - 5%)
(Cost of equity - 5%)=1.05/40       
Cost of equity=1.05/40+5%=0.07625

Total amount of 500 coupon bonds outstanding at $950=500*950=475000
Total amount of 20000 shares outstanding $40=20000*40=800000
Total amount=475000+800000=1275000
Weight of debt=475000/1275000=0.37254902
Weight of equity=800000/1275000=0.62745098

WACC=(Weight of equity)*(Cost of equity)+(Weight of debt)*(Cost of debt)*(1-Tax rate)
WACC=(0.62745098)*(0.07625)+(0.37254902)*(4.64%)*(1-30%)
=4.78%+1.21%
=5.99%

If you have any doubt regarding the solution, please me know.